jason is 1 year older than twice his bro's age. the difference of their ages is 18. find their present ages.
Answer by Paul(988)

LEt the brother be x
Let jason be 2x+1
Difference = 18
(x)-(2x+1)=18
x+1-1=18-1
x=17
Jason = 35 Therefore Jason is 35 and his brother is 17